A car is traveling a road that makes quarter of a
Answer #1

here,

Speed (S) = distance/ time.

In here, distance of one fourth circle = circumf. of circle/2 = Pi*r /2

(displacement , d = sqrt(r^2 + r^2) = 1.41 * r)

Avg. Velocity = displacement/ time

In here, displacement is diameter = d.

therefore,

S = (Pi * r/2)/ t

2*S/Pi = r / t

and

Avg.V = D/t = 1.41 * r/t = 1.41 *2 * S/Pi

Avg V = 0.9 * S